How to solve 7n-2=6n-2n

Answer by ewatrrr(24785)

 
Hi
Idea is to Isolate the variable on one side of the EQ so as You can read the answer
7n-2=6n-2n |Combine 6n-2n = 4n
7n - 2 = 4n |Subtract 4n and Add 2 to both sides of EQ
3n = 2
n = 2/3